# LLM4Drive: A Survey of Large Language Models for Autonomous Driving

### v3

@@ Abstract
Autonomous driving technology, a catalyst for revolutionizing transportation and urban mobility, has the tend to transition from rule-based systems to data-driven strategies. Traditional module-based systems are constrained by cumulative errors among cascaded modules and inflexible pre-set rules. In contrast, end-to-end autonomous driving systems have the potential to avoid error accumulation due to their fully data-driven training process, although they often lack transparency due to their “black box” nature, complicating the validation and traceability of decisions. Recently, large language models (LLMs) have demonstrated abilities including understanding context, logical reasoning, and generating answers. A natural thought is to utilize these abilities to empower autonomous driving. By combining LLM with foundation vision models, it could open the door to open-world understanding, reasoning, and few-shot learning, which current autonomous driving systems are lacking. In this paper, we systematically review a research line about Large Language Models for Autonomous Driving (LLM4AD). This study evaluates the current state of technological advancements, distinctly outlining the principal challenges and prospective directions for the field. For the convenience of researchers in academia and industry, we provide real-time updates on the latest advances in the field as well as relevant open-source resources via the designated link: https://github.com/Thinklab-SJTU/Awesome-LLM4AD.
@@ 1 Introduction
Autonomous driving is rapidly reshaping our understanding of transportation, heralding a new era of technological revolution. This transformation means not only the future of transportation but also a fundamental shift across various industries. In conventional autonomous driving systems, algorithms typically adopt the modular design Liang et al. (2020); Luo et al. (2018); Sadat et al. (2020), with separate components responsible for critical tasks such as perception Li et al. (2022c); Liu et al. (2023d), prediction Shi et al. (2022); Jia et al. (2022a, 2023c), and planning Treiber et al. (2000); Dauner et al. (2023). Specifically, the perception component handles object detection Li et al. (2022c); Liu et al. (2023d), tracking Zeng et al. (2022), and sophisticated semantic segmentation tasks Cheng et al. (2022). The prediction component analyzes the external environment Jia et al. (2021) and estimates the future states of the surrounding agents Jia et al. (2022b). The planning component, often reliant on rule-based decision algorithms Treiber et al. (2000), determines the optimal and safest route to a predetermined destination. While the module-based approach provides reliability and enhanced security in a variety of scenarios, it also presents challenges. The decoupled design between system components may lead to key information loss during transitions and potentially redundant computation as well. Additionally, errors may accumulate within the system due to inconsistencies in optimization objectives among the modules, affecting the vehicle’s overall decision-making performance Chen et al. (2023a).
@@ 2 Motivation of LLM4AD
Tanahashi et al. (2023) ] ] [Question Answering(QA), fill=orange!70 [Traditional QA, fill=orange!70 [Domain Knowledge Distillation Tang et al. (2023)
Human-Centric Autonomous Systems Yang et al. (2023b)] ] [Visual QA, fill=orange!70 [DriveMLM Wang et al. (2023c)
@@
DriveGPT4 Xu et al. (2023)] ] ] [Generation(Diffusion), fill=yellow!70, l sep+=1em [ADriver-I Jia et al. (2023a)
@@
Driving into the Future Wang et al. (2023f)] ] [Planning & Control, fill=blue!30, name=rightNode [Fine-tuning Pre-trained Model, fill=blue!30 [DriveMLM Wang et al. (2023c)
@@
MTD-GPT Liu et al. (2023b)] ] [Prompt Engineering, fill=blue!30 [A Safety Perspective Wang et al. (2023e)
@@
ChatGPT as Your Vehicle Co-Pilot Wang et al. (2023b)
@@
DiLu Wen et al. (2023a)] ] ] [Perception, fill=green!50 [Prediction, fill=green!50 [Can you text what is happening Keysan et al. (2023)
MTD-GPT Liu et al. (2023b)] ] [Detection, fill=green!50 [HiLM-D Ding et al. (2023)] ] [Tracking, fill=green!50 [LanguagePrompt Wu et al. (2023)] ] ] ] z
@@ 3.1.1 Fine-tuning pre-trained models
In the application of fine-tuning pre-trained models, MTD-GPT  Liu et al. (2023b) translates multi-task decision-making problems into sequence modeling problems. Through training on a mixed multi-task dataset, it addresses various decision-making tasks at unsignaled intersections. Although this approach outperforms the performance of single-task decision-making RL models, the used scenes are limited to unsignaled intersections, which might be enough to demonstrate the complexity of the real world application. Driving with LLMs  Chen et al. (2023b) designs an architecture that fuses vectorized inputs into LLMs with a two-stage pretraining and fine-tuning method. Due to the limitation of vectorized representations, their method are only tested in the simulation. DriveGPT4 Xu et al. (2023) presents a multimodal LLM based on Valley Luo et al. (2023a) and develops a visual instruction tuning dataset for interpretable autonomous driving. Besides predicting a vehicle’s basic control signals, it also responds in real-time, explaining why the action is taken. It outperforms baseline models in a variety of QA tasks while the experiments about planning is simple. GPT-Driver Mao et al. (2023a) transforms the motion planning task into a language modeling problem. It exceeds the UniADHu et al. (2023b) in the L2 metric. Nevertheless, since it uses past speed and acceleration information, there is concern about unfair comparison with UniAD. Additionally, L2 only reflects the fitting degree of the driving route and might not reflect the driving performance Dauner et al. (2023). Agent-Driver Mao et al. (2023b) leverages LLMs common sense and robust reasoning capabilities to improve the capabilities of planning by designing a tool library, a cognitive memory, and a reasoning engine. This paradigm achieves better results on the nuScenes dataset. Meanwhile, shortening the inference time is also an urgent problem. DriveLM Sima et al. (2023) uses a trajectory tokenizer to process ego-trajectory signals to texts, making them belong to the same domain space. Such a tokenizer can be applied to any general vision language models. Moreover, they utilize a graph-structure inference with multiple QA pairs in logical order, thus improving the final planning performance. Wang et al. (2023b) adapts LLMs as a vehicle ”Co-Pilot” of driving, which can accomplish specific driving tasks with human intention satisfied based on the information provided. It lacks verification in complex interaction scenarios. LMDrive Shao et al. (2023) designs a multi-modal framework to predict the control signal and whether the given instruction is completed. It adopts Resnet He et al. (2016) as the vision encoder which has not been through an image-text alignment pretraining. In addition, it introduces a benchmark LangAuto which includes approximately 64K instruction-following data clips in CARLA. The LangAuto benchmark tests the system’s ability to handle complex instructions and challenging driving scenario. DriveMLM Wang et al. (2023c) adopts a multi-modal LLM(Multi-view image, Point cloud, and prompt) to generate high-level decision commands and uses Apollo as a planner to get the control signal. Moreover, the training data generated by experts and uses GPT-3.5 to increase data diversity. It achieves 76.1 driving score on the CARLA Town05 Long, which reaches the level of classic end-to-end autonomous driving.
@@
In the prompt engineering perspective, some methods tried to tap into the deep reasoning potential of the LLMs through clever prompt design. DiLu Wen et al. (2023a) designs a framework of LLMs as agents to solve closed-loop driving tasks. This method introduces a memory module to record experience, to leverage LLMs to facilitate reasoning and reflection processes. DiLu exhibits strong generalization capabilities compared with SOTA RL-based methods. However, the reasoning and reflection processes require multiple rounds of question-answering, and its inference time cannot be ignored. Similarly, Receive Reason and React  Cui et al. (2023b) and Drive as You Speak Cui et al. (2023a) integrate the language and reasoning capabilities of LLMs into autonomous vehicles. In addition to memory and reflection processes, these methods introduce additional raw sensor information such as camera, GNSS, lidar, and radar. However, the inference speed is unsolved as well. Furthermore, SurrealDriver Jin et al. (2023b) divides the memory module into short-term memory, long-term guidelines, and safety criteria. Meanwhile, it interviews 24 drivers and uses their detailed descriptions of driving behaviors as chain-of-thought prompts to develop a ‘coach agent’ module. However, there is a lack of comparison with traditional algorithms to prove that large language models indeed bring performance improvements. LanguageMPC Sha et al. (2023) also designs a chain-of-thought framework for LLMs in driving scenarios and it integrates with low-level controllers by guided parameter matrix adaptation. Although its performance exceeds MPC and RL-based methods in the simplified simulator environments, it lacks validation in complex environments. TrafficGPT Zhang et al. (2023b) is a fusion of ChatGPT and traffic foundation models which can tackle complex traffic-related problems and provide insightful suggestions. It leverages multimodal data as a data source, offering comprehensive support for various traffic-related tasks. Talk2BEV Dewangan et al. (2023) introduces a large vision-language model (LVLM) interface for bird’s-eye view (BEV) maps in autonomous driving contexts. It does not require any training or fine-tuning, only relying on pre-trained image-language models. In addition, it presents a benchmark for evaluating subsequent work in LVLMs for AD applications. Talk2Drive Cui et al. (2023a) utilizes human verbal commands and makes autonomous driving decisions based on contextual information to meet humanly personalized preferences for safety, efficiency, and comfort.
@@
MTD-GPT  Liu et al. (2023b) uses single-subtask success rates as the metric in simulation and it exceeds RL expert. DriveGPT4 Xu et al. (2023) uses root mean squared error (RMSE) and threshold accuracies for evaluation. In vehicle action description, justification, and full sentences, it uses BLEU-4 Papineni et al. (2002), METEOR Banerjee and Lavie (2005), CIDERVedantam et al. (2015) and chatgpt score Fu et al. (2023b). Driving with LLMs  Chen et al. (2023b) uses the Mean Absolute Error (MAE) for the predictions of the number of cars and pedestrians, normalized acceleration, and brake pressure. Additionally, it measures the accuracy of traffic light detection as well as the mean absolute distance error in meters for traffic light distance prediction. Besides perception-related metrics, it also uses GPT-3.5 to grade their model’s answers which is a recently emerging technique - grading natural language responses  Fu et al. (2023b); Wang et al. (2023a); Liu et al. (2023c). DiLu Wen et al. (2023a) uses Success Steps in simulation as a metric to evaluate generalization and transformation abilities. SurrealDriver Jin et al. (2023b) evaluates agents based on two main dimensions: safe driving ability and humanness. Safe driving capabilities are assessed through collision rates, while human likeness is assessed through user experiments with 24 adult participants (age 29.3 ± 4.9 years, male = 17 years) who are legal to drive. LanguageMPC Sha et al. (2023) customs some metrics: failure/collision cases, the efficiency of traffic flow, time cost by ego vehicle, and the safety of the ego vehicle’s driving behavior. Similarly, Talk2BEV Dewangan et al. (2023) measures their methods from the perspective of spatial reasoning, instance attribute, instance counting, and visual reasoning. GPT-Driver Mao et al. (2023a), DriveLM Sima et al. (2023) and Agent-Driver Mao et al. (2023b) contains two metrics: L2 error (in meters) and collision rate (in percentage). The average L2 error is calculated by measuring the distance of each waypoint in the planned trajectory and the offline recorded human driver trajectory. It reflects the fitting of the planned trajectory to the human driving trajectory. The collision rate is calculated by placing an ego vehicle box at each waypoint of the planned trajectory and then checking for collisions with the ground truth bounding boxes of other objects. It reflects the safety of the planned trajectory. LMDrive Shao et al. (2023) and DriveMLMWang et al. (2023c) adopt CARLA’s official metrics including Driving Score(DS), Route Completion(RC), Infraction Score(IS). At present, LLM4AD regarding the planning task lacks a unified metric and cannot uniformly evaluate the pros and cons between each method and traditional counterparts.

### v4

@@ Abstract
Autonomous driving technology, a catalyst for revolutionizing transportation and urban mobility, has the tend to transition from rule-based systems to data-driven strategies. Traditional module-based systems are constrained by cumulative errors among cascaded modules and inflexible pre-set rules. In contrast, end-to-end autonomous driving systems have the potential to avoid error accumulation due to their fully data-driven training process, although they often lack transparency due to their ”black box” nature, complicating the validation and traceability of decisions. Recently, large language models (LLMs) have demonstrated abilities including understanding context, logical reasoning, and generating answers. A natural thought is to utilize these abilities to empower autonomous driving. By combining LLM with foundation vision models, it could open the door to open-world understanding, reasoning, and few-shot learning, which current autonomous driving systems are lacking. In this paper, we systematically review the research line about (Vision) Large Language Models for Autonomous Driving ((V)LLM4Drive). This study evaluates the current state of technological advancements, distinctly outlining the principal challenges and prospective directions for the field. For the convenience of researchers in academia and industry, we provide real-time updates on the latest advances in the field as well as relevant open-source resources via the designated link: https://github.com/Thinklab-SJTU/Awesome-LLM4AD.
@@ 1 Introduction
Autonomous driving is rapidly reshaping our understanding of transportation, heralding a new era of technological revolution. This transformation means not only the future of transportation but also a fundamental shift across various industries. In conventional autonomous driving systems, algorithms typically adopt the modular design Liang et al. (2020); Luo et al. (2018); Sadat et al. (2020), with separate components responsible for critical tasks such as perception Li et al. (2022c); Liu et al. (2023d), prediction Shi et al. (2022); Jia et al. (2022a, 2023c, 2024a), and planning Treiber et al. (2000); Dauner et al. (2023); Li et al. (2024b); Jia et al. (2024b). Specifically, the perception component handles object detection Li et al. (2022c); Liu et al. (2023d), tracking Zeng et al. (2022), and sophisticated semantic segmentation tasks Cheng et al. (2022). The prediction component analyzes the external environment Jia et al. (2021) and estimates the future states of the surrounding agents Jia et al. (2022b). The planning component, often reliant on rule-based decision algorithms Treiber et al. (2000), determines the optimal and safest route to a predetermined destination. While the module-based approach provides reliability and enhanced security in a variety of scenarios, it also presents challenges. The decoupled design between system components may lead to key information loss during transitions and potentially redundant computation as well. Additionally, errors may accumulate within the system due to inconsistencies in optimization objectives among the modules, affecting the vehicle’s overall decision-making performance Chen et al. (2023a).
@@ 2 Motivation of LLM4AD
Evaluation of LLMs Tanahashi et al. (2023)
Testing LLMs Tang et al. (2024)
DriveSim Sreeram et al. (2024)
ELM Zhou et al. (2024)
LimSim++ Fu et al. (2024)
OmniDrive Wang et al. (2024d)
AIDE Liang et al. (2024) ] ] [Question Answering(QA), fill=orange!70 [Traditional QA, fill=orange!70 [Domain Knowledge Distillation Tang et al. (2023)
Human-Centric Autonomous Systems Yang et al. (2023b)
Engineering Safety Nouri et al. (2024)
Hybrid Reasoning Azarafza et al. (2024) ] ] [Visual QA, fill=orange!70 [DriveMLM Wang et al. (2023d)
@@
DriveGPT4 Xu et al. (2023)
A Superalignment Framework Kong et al. (2024)
EM-VLM4AD Gopalkrishnan et al. (2024)
TransGPT Wang et al. (2024c) ] ] ] [Generation(Diffusion), fill=yellow!70, l sep+=1em [ADriver-I Jia et al. (2023a)
@@
Driving into the Future Wang et al. (2023g)
ChatScene Zhang et al. (2024)
REvolve Hazra et al. (2024)
GenAD Yang et al. (2024)
DriveDreamer-2 Zhao et al. (2024)
ChatSim Wei et al. (2024)
LLM-Assisted Light Wang et al. (2024a)
LangProp Ishida et al. (2024) ] ] [Planning & Control, fill=blue!30, name=rightNode [Fine-tuning Pre-trained Model, fill=blue!30 [DriveMLM Wang et al. (2023d)
@@
MTD-GPT Liu et al. (2023b)
KoMA Jiang et al. (2024)
AsyncDriver Chen et al. (2024)
PlanAgent Zheng et al. (2024b)
AgentsCoDriver Hu et al. (2024)
DriveVLM Tian et al. (2024)
RAG-Driver Yuan et al. (2024)
VLP Pan et al. (2024)
DME-Driver Han et al. (2024)
] ] [Prompt Engineering, fill=blue!30 [A Safety Perspective Wang et al. (2023f)
@@
ChatGPT as Your Vehicle Co-Pilot Wang et al. (2023c)
@@
DiLu Wen et al. (2023a)
LLM-assited light Wang et al. (2024b)
AccidentGPT Wang et al. (2023b)
LLM-Assist Sharan et al. (2023)
LLaDA Li et al. (2024a) ] ] ] [Perception, fill=green!50 [Prediction, fill=green!50 [Can you text what is happening Keysan et al. (2023)
MTD-GPT Liu et al. (2023b)
LC-LLM Peng et al. (2024)
LeGo-Drive Paul et al. (2024)
Context-aware Motion Prediction Zheng et al. (2024a) ] ] [Detection, fill=green!50 [HiLM-D Ding et al. (2023) ] ] [Tracking, fill=green!50 [LanguagePrompt Wu et al. (2023)] ] ] ]
@@ 3.1.1 Fine-tuning pre-trained models
In the application of fine-tuning pre-trained models, MTD-GPT  Liu et al. (2023b) translates multi-task decision-making problems into sequence modeling problems. Through training on a mixed multi-task dataset, it addresses various decision-making tasks at unsignaled intersections. Although this approach outperforms the performance of single-task decision-making RL models, the used scenes are limited to unsignaled intersections, which might be enough to demonstrate the complexity of the real world application. Driving with LLMs  Chen et al. (2023b) designs an architecture that fuses vectorized inputs into LLMs with a two-stage pretraining and fine-tuning method. Due to the limitation of vectorized representations, their method are only tested in the simulation. DriveGPT4 Xu et al. (2023) presents a multimodal LLM based on Valley Luo et al. (2023a) and develops a visual instruction tuning dataset for interpretable autonomous driving. Besides predicting a vehicle’s basic control signals, it also responds in real-time, explaining why the action is taken. It outperforms baseline models in a variety of QA tasks while the experiments about planning is simple. GPT-Driver Mao et al. (2023a) transforms the motion planning task into a language modeling problem. It exceeds the UniADHu et al. (2023b) in the L2 metric. Nevertheless, since it uses past speed and acceleration information, there is concern about unfair comparison with UniAD. Additionally, L2 only reflects the fitting degree of the driving route and might not reflect the driving performance Dauner et al. (2023). Agent-Driver Mao et al. (2023b) leverages LLMs common sense and robust reasoning capabilities to improve the capabilities of planning by designing a tool library, a cognitive memory, and a reasoning engine. This paradigm achieves better results on the nuScenes dataset. Meanwhile, shortening the inference time is also an urgent problem. DriveLM Sima et al. (2023) uses a trajectory tokenizer to process ego-trajectory signals to texts, making them belong to the same domain space. Such a tokenizer can be applied to any general vision language models. Moreover, they utilize a graph-structure inference with multiple QA pairs in logical order, thus improving the final planning performance. Wang et al. (2023c) adapts LLMs as a vehicle ”Co-Pilot” of driving, which can accomplish specific driving tasks with human intention satisfied based on the information provided. It lacks verification in complex interaction scenarios. LMDrive Shao et al. (2023) designs a multi-modal framework to predict the control signal and whether the given instruction is completed. It adopts Resnet He et al. (2016) as the vision encoder which has not been through an image-text alignment pretraining. In addition, it introduces a benchmark LangAuto which includes approximately 64K instruction-following data clips in CARLA. The LangAuto benchmark tests the system’s ability to handle complex instructions and challenging driving scenario. DriveMLM Wang et al. (2023d) adopts a multi-modal LLM(Multi-view image, Point cloud, and prompt) to generate high-level decision commands and uses Apollo as a planner to get the control signal. Moreover, the training data generated by experts and uses GPT-3.5 to increase data diversity. It achieves 76.1 driving score on the CARLA Town05 Long, which reaches the level of classic end-to-end autonomous driving. KoMA Jiang et al. (2024) is a knowledge-driven multi-agent framework in which each agent is powered by large language models. These agents analyze and infer the intentions of surrounding vehicles to enhance decision-making. AsyncDriver Chen et al. (2024) is an asynchronous LLM-enhanced framework where the inference frequency of LLM is controllable and can be decoupled from the real-time planner. It has good closed-loop evaluation performance in challenging scenarios of nuPlan. PlanAgent Zheng et al. (2024b) extracts bird’s-eye view (BEV) representation and generates a text description input based on the lane map through an environment transformation module. It uses a reasoning engine module to perform a hierarchical thinking chain to guide driving scene understanding, motion command generation, and planning code writing. AGENTSCODRIVER Hu et al. (2024) is an LLM-powered framework for multi-vehicle collaborative driving with lifelong learning, enabling communication and collaboration among driving agents in complex traffic scenarios, featuring a reasoning engine, cognitive memory, reinforcement reflection, and a communication module. DriveVLM Tian et al. (2024) leverages Vision-Language Models to enhance scene understanding and planning capabilities for autonomous driving, while DriveVLM-Dual synergizes these advancements with traditional 3D perception and planning approaches to effectively address spatial reasoning and computational challenges, demonstrating superior performance in complex and dynamic driving scenarios. RAG-Driver Yuan et al. (2024), a Multi-Modal Large Language Model with Retrieval-augmented In-context Learning, provides explainable and generalizable end-to-end driving by producing numerical control signals, along with explanations and justifications for driving actions, and demonstrates impressive zero-shot generalization to unseen environments without additional training. LLaDA Li et al. (2024a) designs a training-free mechanism to assist human drivers and adapt autonomous driving policies to new environments. VLP Pan et al. (2024) is a Vision-Language-Planning model intended to enhance autonomous driving systems (ADS) by incorporating two novel components: ALP and SLP. ALP (Agent-wise Learning Paradigm) aligns the generated bird’s-eye-view (BEV) with the true BEV map, improving self-driving BEV reasoning. SLP (Self-Driving-Car-Centric Learning Paradigm) aligns the ego vehicle’s query features with its textual planning features, enhancing self-driving decision-making. DME-Driver Han et al. (2024) enhances decision logic explainability and environmental perception accuracy by using a vision language model for decision-making and a planning-oriented perception model for generating precise control signals, effectively translating human-like driving logic into actionable commands, and achieving high-precision planning accuracy through the comprehensive HBD dataset.
@@
In the prompt engineering perspective, some methods tried to tap into the deep reasoning potential of the LLMs through clever prompt design. DiLu Wen et al. (2023a) designs a framework of LLMs as agents to solve closed-loop driving tasks. This method introduces a memory module to record experience, to leverage LLMs to facilitate reasoning and reflection processes. DiLu exhibits strong generalization capabilities compared with SOTA RL-based methods. However, the reasoning and reflection processes require multiple rounds of question-answering, and its inference time cannot be ignored. Similarly, Receive Reason and React  Cui et al. (2023b) and Drive as You Speak Cui et al. (2023a) integrate the language and reasoning capabilities of LLMs into autonomous vehicles. In addition to memory and reflection processes, these methods introduce additional raw sensor information such as camera, GNSS, lidar, and radar. However, the inference speed is unsolved as well. Furthermore, SurrealDriver Jin et al. (2023b) divides the memory module into short-term memory, long-term guidelines, and safety criteria. Meanwhile, it interviews 24 drivers and uses their detailed descriptions of driving behaviors as chain-of-thought prompts to develop a ‘coach agent’ module. However, there is a lack of comparison with traditional algorithms to prove that large language models indeed bring performance improvements. LanguageMPC Sha et al. (2023) also designs a chain-of-thought framework for LLMs in driving scenarios and it integrates with low-level controllers by guided parameter matrix adaptation. Although its performance exceeds MPC and RL-based methods in the simplified simulator environments, it lacks validation in complex environments. TrafficGPT Zhang et al. (2023b) is a fusion of ChatGPT and traffic foundation models which can tackle complex traffic-related problems and provide insightful suggestions. It leverages multimodal data as a data source, offering comprehensive support for various traffic-related tasks. Talk2BEV Dewangan et al. (2023) introduces a large vision-language model (LVLM) interface for bird’s-eye view (BEV) maps in autonomous driving contexts. It does not require any training or fine-tuning, only relying on pre-trained image-language models. In addition, it presents a benchmark for evaluating subsequent work in LVLMs for AD applications. Talk2Drive Cui et al. (2023a) utilizes human verbal commands and makes autonomous driving decisions based on contextual information to meet humanly personalized preferences for safety, efficiency, and comfort. AccidentGPT  Wang et al. (2023b) integrates multi-vehicle collaborative perception to improve environmental understanding and collision avoidance, offering advanced safety features like proactive remote safety warnings and blind spot alerts. It also supports traffic police and management agencies by providing real-time intelligent analysis of traffic safety factors.
@@
MTD-GPT  Liu et al. (2023b) uses single-subtask success rates as the metric in simulation and it exceeds RL expert. DriveGPT4 Xu et al. (2023) and RAG-Driver Yuan et al. (2024) uses root mean squared error (RMSE) and threshold accuracies for evaluation. In vehicle action description, justification, and full sentences, it uses BLEU-4 Papineni et al. (2002), METEOR Banerjee and Lavie (2005), CIDERVedantam et al. (2015) and chatgpt score Fu et al. (2023b). Driving with LLMs  Chen et al. (2023b) uses the Mean Absolute Error (MAE) for the predictions of the number of cars and pedestrians, normalized acceleration, and brake pressure. Additionally, it measures the accuracy of traffic light detection as well as the mean absolute distance error in meters for traffic light distance prediction. Besides perception-related metrics, it also uses GPT-3.5 to grade their model’s answers which is a recently emerging technique - grading natural language responses  Fu et al. (2023b); Wang et al. (2023a); Liu et al. (2023c). DiLu Wen et al. (2023a) uses Success Steps in simulation as a metric to evaluate generalization and transformation abilities. SurrealDriver Jin et al. (2023b) evaluates agents based on two main dimensions: safe driving ability and humanness. Safe driving capabilities are assessed through collision rates, while human likeness is assessed through user experiments with 24 adult participants (age 29.3 ± 4.9 years, male = 17 years) who are legal to drive. LanguageMPC Sha et al. (2023) customs some metrics: failure/collision cases, the efficiency of traffic flow, time cost by ego vehicle, and the safety of the ego vehicle’s driving behavior. Similarly, Talk2BEV Dewangan et al. (2023) measures their methods from the perspective of spatial reasoning, instance attribute, instance counting, and visual reasoning. GPT-Driver Mao et al. (2023a), LLaDA Li et al. (2024a), DriveLM Sima et al. (2023), Agent-Driver Mao et al. (2023b), VLP Pan et al. (2024), DME-Driver Han et al. (2024) and DriveVLM Tian et al. (2024) contain two metrics: L2 error (in meters) and collision rate (in percentage). The average L2 error is calculated by measuring the distance of each waypoint in the planned trajectory and the offline recorded human driver trajectory. It reflects the fitting of the planned trajectory to the human driving trajectory. The collision rate is calculated by placing an ego vehicle box at each waypoint of the planned trajectory and then checking for collisions with the ground truth bounding boxes of other objects. It reflects the safety of the planned trajectory. LMDrive Shao et al. (2023) and DriveMLMWang et al. (2023d) adopts CARLA’s official metrics including Driving Score(DS), Route Completion(RC), Infraction Score(IS). At present, LLM4AD regarding the planning task lacks a unified metric and cannot uniformly evaluate the pros and cons between each method and traditional counterparts. KoMA Jiang et al. (2024) demonstrates the effectiveness and high success rate in the Highway MARL simulator. AsyncDriver Chen et al. (2024) obtains superior closed-loop evaluation performance in nuPlan Closed-Loop Reactive Hard20 scenarios. PlanAgent Zheng et al. (2024b) achieves competitive and general results on the nuPlan Val14 and Test14-hard benchmarks, and improves the efficiency of token usage when describing driving scenarios. AGENTSCODRIVER Hu et al. (2024) adopt Success Rate (SR) and Success Step (SS) in HighwayEnv simulator Leurent (2018). LeGo-Drive Paul et al. (2024) is a novel planning-guided end-to-end LLM-based goal point navigation solution that predicts and improves the desired state by dynamically interacting with the environment and generating a collision-free trajectory.